OpenStudy (anonymous):

what are 2 possible combinations of nickels and quarters in $3.80?

OpenStudy (anonymous):

15 quarters and 1 nickel (.25*15 =3.75) + 1 nickel(.05) = 3.80 and 12 quarters (12*.25) = 3.00 + 16 nickels (.05*16) = 3.80